OpenStudy (maya):

∑from 1to infinity of n^2* e^-2n

OpenStudy (maya):

Does this converge or diverge?

OpenStudy (anonymous):

converges like crazy

OpenStudy (maya):

converges to?

OpenStudy (anonymous):

\[\sum_{k=1}^\infty \frac{n^2}{e^{2n}}\]

OpenStudy (anonymous):

who knows. just that the denominator grows tons faster than numerator. i will cheat and find the sum if you like

OpenStudy (maya):

only if you can tell me how you got it? please

OpenStudy (anonymous):

i have absolutely no idea how to add this up.

OpenStudy (anonymous):

leave the question open and someone may know. i do know that the answer is \[\frac{e^2(e^2+1)}{(e^2-1)^3}\]

OpenStudy (maya):

alright thanks

OpenStudy (anonymous):

but no idea how to sum. keep asking

OpenStudy (anonymous):

hey maya I think I know how to do this

OpenStudy (anonymous):

give me a bit to work it out

OpenStudy (anonymous):

ok I have the solution let me type it out:

OpenStudy (maya):

great

OpenStudy (anonymous):

let w=e^-2. Then our sum is: \[\sum_{n=0}^{\infty}n^2w^n\] now consider the following sum: \[\sum_{n=0}^{\infty}w^n\] we will take its second derivitive: \[\left( \sum_{n=0}^{\infty} w^n\right) ''=\sum_{n=0}^{\infty}n*(n-1)*w^n=\sum_{n=0}^{\infty}n^2w^{n-2}-\sum_{n=0}^{\infty}nw^{n-2}\] we know the geometric series' solution on the left so we can find its second derivative: \[\sum_{n=0}^{\infty}w^n=\frac{1}{1-w}\] \[\left( \frac{1}{1-w} \right)''=\frac{2}{(1-w)^3}\] multiplying both sides by w^2 \[\frac{2w^2}{(1-w)^3}=\sum_{n=0}^{\infty}n^2w^n-\sum_{n=0}^{\infty}nw^n\]

OpenStudy (anonymous):

ok almost there, now we have the sum you want in the equation

OpenStudy (anonymous):

so we use the same trick for the rightmost sum: \[\frac{w}{(1-w)^2}=\sum_{n=0}^{\infty}nw^n\] now we plug in to find the sum we wanted: \[\frac{2w^2}{(1-w)^3}+\frac{w}{(1-w)^2}=\sum_{n=0}^{\infty}n^2w^n\] plug in e^-2=w and simplify and you get the answer satellite posted

OpenStudy (anonymous):

ok so that is a lot to take in probably let me know if you have a question

OpenStudy (anonymous):

comes out to: \[\frac{e^2(1+e^2)}{(e^2-1)^3}\]

OpenStudy (maya):

hmm trying this myself. I think i get it

OpenStudy (maya):

thank you

OpenStudy (anonymous):

you're welcome :)

OpenStudy (maya):

prettyyy confusing would have never thought of doing that myself

OpenStudy (anonymous):

we did stuff like this a lot in complex analysis that's the only reason I knew to do it.

OpenStudy (anonymous):

oh small mistake when taking the second derivative of geometric series should be n*(n-1)*w^(n-2) not just w^n

OpenStudy (anonymous):

just a typo

OpenStudy (maya):

ok

OpenStudy (anonymous):

WOW! what gave you the idea to take the second derivative??

OpenStudy (anonymous):

trying to get the n^2 term in front :)

OpenStudy (anonymous):

aha! very nice

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

actually very very nice

OpenStudy (anonymous):

my complex analysis prof liked doing similar probs

OpenStudy (anonymous):

really good prof, hes the one doing galois theory in the fall

OpenStudy (anonymous):

then you will have lots of fun:)